# FedRTS: Federated Robust Pruning via Combinatorial Thompson Sampling

### v1

@@ C.1 Greedy Optimal Action
$\begin{split}&\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{B_{t}}\Delta_{S_{t}}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{B_{t}}\right]\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{|\mu_{i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}}\right]\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\land|\mu_{i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\right)\Delta_{\max}\\ &=\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\\ &\cdot\mathbb{P}\left(|\mu_{i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\middle|N_{t,i}=\tau\right)\Delta_{\max}.\end{split}$ (28)
@@
$\begin{split}&\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\\ &\cdot\mathbb{P}\left(|\mu_{i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\middle|N_{t,i}=\tau\right)\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\cdot 2\exp(-3\log T)\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\frac{2}{T}\Delta_{\max}\\ &=2\Delta_{\max}\langle W\rangle.\end{split}$ (29)
@@
$\begin{split}\mathbb{E}&\left[\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{\neg A_{t}\land\neg B_{t}\land C_{t,k}}\right]\\ &\leq\mathbb{E}\left[\sum_{s\notin S^{*}_{k}}\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{s=s_{t,k}\land N_{t,s}\leq F(s)}\right]+\frac{U\Delta_{max}}{\epsilon^{6}}.\end{split}$ (30)
Proof. This Lemma is a special case in (Kong et al., 2021) (Lemma 1), setting the size of unit as 1 and $U=CC^{\prime}$.
@@
$\begin{split}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{s^{*}_{k}=s_{t,k}\land||\xi_{t,s^{*}_{k}}-\mu_{s^{*}_{k}}||_{\infty}\leq\epsilon}\right]\Delta_{\max}\leq 2+\frac{8}{\epsilon^{2}}\Delta_{\max}\end{split}$ (31)
Proof. This Lemma is a special case in (Kong et al., 2021) (Lemma 3), setting the size of each unit as 1.
@@
Proof. This Lemma has been proved in (Wang & Chen, 2018) (Lemma 3).
@@
Proof. This Lemma can be proved by Chernorff-Hoeffding Bound in (Phillips, 2012)(Theorem 1.1).
Appendix D Additional Setup Details
This section provides supplementary information to the main contents, elaborating on the datasets, hyperparameter details, and baseline frameworks used in our experiments.
D.1 Datasets
In the CV tasks, we apply our approach using ResNet18 on CIFAR-10, CINIC-10 and SVHN datasets.
D.2 Hyperparameter Details
In addition to the Federated Learning settings described earlier, we provide more detailed hyperparameter configurations. In the TSAdj, we define $\kappa=0.2\left(4-\cos{\frac{t\pi}{T_{end}}}\right)K^{l}$ at $t$-th loop, where $K^{l}$ represents the number of active parameters in $l$-th layer. The optimizer used is Adam, with a learning rate of $\eta=0.001$. To generate non-iid data distributions on the client, we utilize the Dirichlet distribution. For CV tasks, we set the heterogeneity degree $a$ to 0.5, which controls the distribution of labels among local datasets. In the NLP tasks, we apply the Dirichlet distribution to create data imbalance, setting $a$ to 5.
D.3 Baselines
  • FedTiny: Initializes the model by adaptive batch normalization selection and adjusts the model mask based on the aggregated top-K gradients of parameters. We disable the adaptive batch normalization selection due to high communication overhead.
  • FedMef: Utilizes budget-aware extrusion to transfer essential information of pruned parameters to other parameters and reduces activation memory by scaled activation pruning. In our experiment, we disable the scaled activation pruning due to computational overhead.
Appendix E Additional Experiments Results
@@
E.1 Impact of trade-off ratio $\gamma$
The trade-off ratio $\gamma$ modulates the weight of individual models’ information. Traditional methods only incorporate the information from the aggregated model, which is insufficient. To verify the impact of individual models’ information, we conduct experiments on the trade-off ratio $\gamma$. As illustrated in Fig. 8, the test accuracy reaches its peak when the $\gamma$ value is between 0.3 and 0.5. This indicates that the information from both individual models and the aggregated model is significant in model topology adjustment. Consequently, the $\gamma$ value is set to 0.5 in the main experiments.
@@ E.1 Impact of trade-off ratio $\gamma$
E.2 Impact of the scaling factor $\lambda$
The scaling factor $\lambda$ regulates the impact of the current reward $X(t)$ at step $t$. A $\lambda$ value that is too small will amplify the uncertainty in Thompson Sampling, whereas a $\lambda$ value that is too large will disregard the uncertainty arising from sample insufficiency. Thus, selecting an appropriate value for $\lambda$ is crucial for enhancing FedRTS’s performance. In a separate set of experiments, the influence of the reward scaling $\lambda$ on the results is investigated. Fig. 9 demonstrates that the test accuracy attains its maximum at approximately 0.73 when $\lambda$ is 10. This finding indicates that a $\lambda$ value of 10 is a reasonable choice, and therefore, we adopt $\lambda=10$ in our main experiments to ensure optimal accuracy.
@@ E.2 Impact of the scaling factor $\lambda$
E.3 The Impact of Adjustment Interval
When considering federated dynamic sparse training methods, the timing of adjustments ($\Delta T$) between rounds holds paramount importance. Striking the right balance in this interval is crucial—a lengthy interval may lead to sparse adjustments, hindering optimal topology discovery, while a short interval may result in inadequate training periods between adjustments. In cases where a method struggles to swiftly recuperate from information loss post-topology changes, it can lead to unreliable adjustments and subpar outcomes. Our evaluation of FedRTS on the CIFAR-10 dataset involved exploring various adjustment intervals ($\Delta T\in\{3,5,10,20\}$). As depicted in Table 3, FedRTS consistently outperformed baseline methods across most intervals. Notably, FedRTS’s performance improved with shorter intervals, showcasing its adeptness in efficiently recuperating from information loss post-topology adjustments—a trait starkly contrasting with FedDST’s diminishing performance under similar conditions.
Table 3: Performance under Different Adjustment Interval ($\Delta T$)
@@ E.3 The Impact of Adjustment Interval
E.4 Convergence
To compare the convergence speed of our proposed FedRTS with other state-of-the-art (SOTA) methods, we conducted experiments on validation accuracy using the CIFAR-10 and CINIC-10 datasets, focusing on cumulative FLOPs and Cumulative Communication Cost during training. On the CIFAR-10 dataset, our proposed FedRTS demonstrated superior accuracy compared to other SOTA frameworks in terms of cumulative FLOPs throughout the entire training period, as depicted in Figure 10. On the CINIC-10 dataset, initially, our proposed FedRTS lagged behind FedDST at the start of training but later outperformed almost all other SOTA frameworks in terms of cumulative communication cost and cumulative FLOPs for the remainder of the training period, as illustrated in Figure 11 and Figure 12. Although FedAVG outperformed FedRTS on CINIC-10 within a certain range of cumulative communication costs (between $150-200GB$), FedRTS surpassed it thereafter, showcasing excellent convergence speed.
In Figure 10, initially, there are no apparent differences in cumulative FLOPs between $0-3\times 10^{15}$, but thereafter, FedRTS consistently outperforms all other methods. Between cumulative FLOPs of 11-13 ×$\times 10^{15}$, FedMef briefly outperforms FedRTS, but beyond that point, FedRTS clearly surpasses all other methods. This pattern showcases the exceptional convergence speed of FedRTS in cumulative FLOPs on CIFAR-10.
In Figure 11, FedDST initially outperforms FedRTS and other methods within the range of cumulative communication costs between $10-50GB$. However, FedRTS overtakes FedDST and other frameworks thereafter. When the cumulative communication cost exceeds $120GB$, the validation accuracy of FedRTS experiences a temporary decline, resulting in lower performance than FedDST. Nonetheless, FedRTS surpasses FedDST once the cumulative communication cost surpasses $160GB$. Notably, FedAVG outperforms all other methods between cumulative communication costs of $150-190GB$ but declines afterward. This phenomenon may stem from the stronger stability of FedAVG, allowing it to converge while other methods begin to fluctuate. Overall, Figure 11 highlights the outstanding convergence speed of FedRTS in cumulative communication costs on CINIC-10.
In Figure 12, FedDST initially surpasses FedRTS and other methods within the range of cumulative FLOPs between $1-15\times 10^{15}$. However, FedRTS overtakes FedDST and outperforms other frameworks thereafter. Although FedDST outperforms FedRTS when cumulative FLOPs range between $35-44\times 10^{15}$, FedRTS surpasses FedDST and other methods beyond that threshold. Overall, Figure 12 demonstrates the excellent convergence speed of FedRTS in cumulative FLOPs on CINIC-10.
@@ E.4 Convergence
Appendix F Communication and Computational Cost

### v3

@@ C.2.1 Lemmas
$\begin{split}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{B_{t}}\Delta_{S_{t}}\right]&\leq\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{B_{t}}\right]\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{|\mu_{i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}}\right]\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\land|\mu_{i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\right)\Delta_{\max}\\ &=\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\cdot\mathbb{P}\left(|\mu_{i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\middle|N_{t,i}=\tau\right)\Delta_{\max}.\end{split}$ (27)
@@
$\begin{split}\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\cdot&\mathbb{P}\left(|\mu_{i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\middle|N_{t,i}=\tau\right)\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\cdot 2\exp(-3\log T)\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\frac{2}{T}\Delta_{\max}\\ &=2\Delta_{\max}\langle W\rangle.\end{split}$ (28)
@@
$\begin{split}\mathbb{E}&\left[\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{\neg A_{t}\land\neg B_{t}\land C_{t,k}}\right]\leq\mathbb{E}\left[\sum_{s\notin S^{*}_{k}}\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{s=s_{t,k}\land N_{t,s}\leq F(s)}\right]+\frac{U\Delta_{max}}{\epsilon^{6}}.\end{split}$ (29)
Proof. This Lemma is a special case in [26] (Lemma 1), setting the size of the unit as 1 and $U=CC^{\prime}$.
@@
[…]
Proof. This Lemma is a special case in [26] (Lemma 3), setting the size of each unit as 1.
@@
Proof. This Lemma has been proved in [59] (Lemma 3).
@@
Proof. This Lemma can be proved by Chernorff-Hoeffding Bound in [44](Theorem 1.1).
Appendix D Limitation
While FedRTS with TSAdj effectively addresses key challenges: greedy adjustments, unstable topologies, and communication inefficiency in federated dynamic pruning, our approach shares a common limitation with prior magnitude-based pruning methods [12, 46]. The theoretical analysis relies on Assumption 3.2, which presumes independence among importance scores. In practice, weight magnitudes exhibit only asymptotic independence [53], creating a minor but acknowledged gap between theory and empirical results.
We considered alternative importance metrics like Fisher information [52] that might better satisfy the independence requirement. However, their computational complexity makes them impractical for federated settings with resource-constrained clients. Following established practice in the pruning literature, we therefore adopt magnitude-based pruning while transparently acknowledging this theoretical-empirical discrepancy, a compromise that reflects the inherent trade-offs in practical federated learning systems.
While the mean-field approximation in Assumption 3.3 provides theoretical grounding for TSAdj, its validity weakens for extremely small models (<100K parameters) where finite-size effects become significant [53]. However, at such small scales, the benefits of pruning become marginal, as these models can typically be trained directly on-device without compression. Thus, this limitation has minimal practical impact on FedRTS.
The FedRTS also presents the following two limitations. Fairness and bias: The method dynamically adjusts model structure per global feedback. In highly heterogeneous settings, this could unintentionally bias the model toward data-rich or majority-class clients. Security/privacy considerations: While the method reduces communication cost, sharing top gradient indices might still leak sensitive information.
Appendix E Broader Impact
Beyond its technical contributions, FedRTS holds significant practical potential for democratizing efficient federated learning. By substantially reducing communication and computation costs via dynamic sparse training, our approach could broaden the adoption of FL in resource-constrained environments, such as the Internet of Things [31, 4, 2, 3, 58], edge perception [61, 62, 60], and privacy-sensitive medical applications [16]. However, like all pruning-based methods, FedRTS carries a risk of inadvertently amplifying biases if sparsity patterns disproportionately discard weights that encode features of minority classes. To mitigate this, we recommend integrating our method with fairness-aware regularization and continual learning techniques [45, 63] in sensitive domains. Future work will explicitly investigate the societal implications of deploying FedRTS at scale, with a focus on applications where model sparsity may impact predictive reliability.
Appendix F Additional Experiments Results
@@
F.1 The Impact of Hyperparameters
F.1.1 The Impact of trade-off ratio $\gamma$
The trade-off ratio $\gamma$ modulates the weight of individual models’ information. Traditional methods only incorporate the information from the aggregated model, which is insufficient. To verify the impact of individual models’ information, we conduct experiments on the trade-off ratio $\gamma$. As illustrated in Fig. 8, the test accuracy reaches its peak when the $\gamma$ value is between 0.3 and 0.5. This indicates that the information from both individual models and the aggregated model is significant in model topology adjustment.
@@ F.1.1 The Impact of trade-off ratio $\gamma$
Table 2: Performance under Different Adjustment Interval ($\Delta T$)
CIFAR-10 CINIC-10
FLOPs Comm. Cost FLOPs Comm. Cost
FedDST 3.4E16 221.9 GB 8.2E16 290.7 GB
FedMef 3.5E16 227.6 GB 6.1E16 222.3 GB
FedRTS 3.1E16 204.2 GB 2.9E16 105.6 GB
Table 3: The overall training FLOPs and communicational cost (Comm. Cost) for FedRTS and other frameworks to reach the accuracy of dense FedAVG.
F.1.2 The Impact of the scaling factor $\lambda$
The scaling factor $\lambda$ regulates the impact of the current reward $X(t)$ at step $t$. A $\lambda$ value that is too small will amplify the uncertainty in Thompson Sampling, whereas a $\lambda$ value that is too large will disregard the uncertainty arising from insufficient samples. Thus, selecting an appropriate value for $\lambda$ is crucial for enhancing FedRTS’s performance. In a separate set of experiments, the influence of the reward scaling $\lambda$ on the results is investigated. Figure 8 shows that the test accuracy attains its maximum at approximately 0.73 when $\lambda$ is 10.
F.1.3 The Impact of the number of core links $\kappa$
The number of core links $\kappa$ critically controls the adjustment intensity (the number of weights pruned and reactivated) in each outer loop, creating an essential trade-off: large $\kappa$ values lead to sluggish topology evolution, while small $\kappa$ risks substantial information loss. We formalize this through our decay schedule: $\kappa^{l}=K^{l}-\frac{\alpha_{adj}}{2}(1+\cos{(\frac{t\pi}{T_{end}})})K^{l}$, where $K^{l}$ is the number of active weights at $l$-th layer. Our empirical investigation on CIFAR-10 in Fig. 14 demonstrates FedRTS’s robustness across varying $\alpha_{adj}$, consistently outperforming baseline methods.
F.1.4 The Impact of Adjustment Interval $\Delta T$
In federated dynamic sparse training, the adjustment interval ($\Delta T$) between two outer loops for adjustment critically influences model performance through a fundamental trade-off. Excessively long intervals may delay the discovery of optimal sparse patterns, while overly frequent adjustments can prevent adequate model adaptation between updates and require higher resource consumption.
Our experiments with FedRTS on CIFAR-10 explored this trade-off by evaluation with various adjustment intervals ($\Delta T\in\{3,5,10,20\}$). As depicted in Table 2, FedRTS consistently outperformed baseline methods across most cases. Notably, FedRTS’s performance improved with shorter intervals, demonstrating FedRTS’s unique ability to quickly recover from topology changes, which is a crucial advantage in dynamic sparse training scenarios.
F.2 Efficiency Analysis
F.2.1 Training Cost
We evaluate FedRTS’s efficiency by analyzing training costs on CIFAR-10 and CINIC-10. To compare expenses, we measure total training FLOPs and communication costs required to match the testing accuracy of dense FedAVG. The accuracy is averaged over 10 rounds to mitigate variance from random factors, ensuring a more stable evaluation.
As shown in Tab. 3, FedRTS achieves the testing accuracy of dense FedAVG with significantly lower costs. On CINIC-10, it achieves 48% training FLOPs and the communication cost compared to the best baseline FedMef, which highlights the efficiency of FedRTS.
F.2.2 Convergence Behavior
Convergence is crucial for effective learning, precise predictions, and optimized resource use. To evaluate this, we track the validation accuracy of FedRTS and SOTA frameworks on CIFAR-10 and CINIC-10 against cumulative communication and computational costs, as shown in Fig. 10, Fig. 10, Fig. 12, and Fig. 12. Accuracy is averaged over 10 rounds to reduce variance and provide a stable representation. Our experimental results demonstrate distinct convergence patterns across datasets. On CIFAR-10, FedRTS exhibits superior convergence characteristics throughout the entire training process. The CINIC-10 dataset reveals a more complex trajectory: while FedDST initially demonstrates higher performance during early training phases, FedRTS consistently surpasses competing methods in later stages when considering both communication efficiency and computational costs. These empirical results provide strong evidence for the effectiveness of the TSAdj mechanism in FedRTS, particularly in optimizing the trade-off between model performance and resource consumption during federated training.
@@ F.2.2 Convergence Behavior
Figure 13: Accuracy on CIFAR-10 with ResNet under various degrees of adjustment ratio $\alpha_{adj}$.
Figure 14: Perplexity on TinyStories under various degrees of data heterogeneity. The lower $a$ represents a higher non-IID degree.
F.3 Broader Experiments
F.3.1 FedRTS vs. FedCUCB
The Combinatorial Upper Confidence Bound (CUCB) is another optimization method for the Combinatorial Multi-Armed Bandit (CMAB) problem. Unlike Combinatorial Thompson Sampling (CTS), which relies on probabilistic sampling, CUCB employs deterministic optimistic reward estimates to balance exploration and exploitation. However, its inherent optimism can lead to persistently selecting suboptimal arm combinations, particularly in dynamic pruning scenarios where adaptability is crucial.
CUCB maintains upper confidence bounds (UCBs) rather than probability distributions for each arm $i$, which formulated as
$\bar{\mu}_{t,i}=\hat{\mu}_{t,i}+\sqrt{\frac{3\log t}{2\Psi_{i}}},$ (33)
where $\Psi_{i}$ denotes the number of times arm $i$ has been selected, $\hat{\mu}_{t,i}$ denotes the empirical mean reward (initialized as $\hat{\mu}_{0,i}=0$). The selected arms for action $S_{t}$ are determined by
$S_{t}=\{i\in\mathrm{Top}(\bar{\mu}_{t},K)\}$
For the selected arms $i\in S_{t}$ with outcomes $X_{t}$,the empirical mean is updated as:
$\Psi_{i}\leftarrow\begin{cases}\Psi_{i},&\text{$i\notin S_{t}$}\\ \Psi_{i}+1,&\text{$i\in S_{t}$}\end{cases},\quad\hat{\mu}_{t+1,i}=\begin{cases}\hat{\mu}_{t,i},&\text{$i\notin S_{t}$}\\ \frac{\hat{\mu}_{t,i}(\Psi_{i}-1)+X_{t,i}}{\Psi_{i}},&\text{$i\in S_{t}$}\end{cases},$ (34)
The next round’s action selection then uses the updated UCB $\bar{\mu}_{t+1,i}$ computed by Eq. 33.
A notable drawback of FedCUCB is its handling of under-explored ("bad") arms. If an arm $i$ is rarely selected (low $\Psi_{i}$), its confidence bound $\bar{\mu}_{t,i}$ grows indefinitely as $t\rightarrow\infty$, eventually forcing its selection. This behavior disrupts convergence, as the algorithm repeatedly revisits suboptimal arms instead of refining the topology.
To evaluate this limitation, we propose FedCUCB, a federated pruning framework based on CUCB. In contrast to FedRTS, FedCUCB maintains UCB $\bar{\mu}_{t}$ and define the outcomes $X_{t}$ as same in Eq. 3. The results in Fig. 15 demonstrates that FedCUCB performs slightly worse than FedRTS, but still outperform other baselines.
Figure 15: Accuracy on CIFAR-10 with ResNet18 for FedRTS, FedCUCB and other federated dynamic pruning methods.
Method Comm. Cost FLOPs Memory Acc.
FedAVG 89.52 MB 17.8E+12 134.29MB 77.24
FedPAQ 22.48 MB 17.8E+12 134.29 MB 73.34
FedDST 20.85 MB 3.8E+12 31.28 MB 68.31
FedRTS 20.85 MB 3.8E+12 31.28 MB 73.41
Table 4: Performance and training cost comparison between federated pruning (FedDST, FedRTS) and federated quantization (FedPAQ) frameworks on CIFAR-10 using ResNet18. FedPAQ apply 8-bit quantization, while FedDST and FedRTS employ a target density of $d^{\prime}=0.2$.
F.3.2 Pruning vs. Weights Quantization
Parallel to pruning-based approaches, quantization techniques [50, 23, 15, 17, 20] offer an alternative for reducing communication overhead in FL. While both strategies aim to alleviate transmission costs, weights quantization operates after dense local training, whereas pruning enables sparse training by design. This fundamental difference grants pruning distinct advantages in resource-constrained settings: it reduces computation through selective updates, lowers memory usage by maintaining dynamic sparsity, and ensures consistent communication efficiency across rounds.
Empirical evidence supports these trade-offs. As shown in Table 4, our FedRTS outperforms 8-bit FedPAQ [50], a weights quantization baseline, in computational and memory efficiency, while retaining competitive model accuracy. This makes pruning a more holistic solution for FL deployments on edge devices.
Appendix G Additional Setup Details
G.1 Datasets
In the CV tasks, we apply our approach using ResNet18 on CIFAR-10, CINIC-10, TinyImageNet, and SVHN datasets.
  • TinyImageNet: It contains 200 classes with 500 training images, 50 validation images, and 50 test images per class, each resized to 64×64 pixels.
G.2 Baselines
  • FedTiny: Initializes the model by adaptive batch normalization selection and adjusts the model mask based on the aggregated top-K gradients of parameters.
  • FedMef: Utilizes budget-aware extrusion to transfer essential information of pruned parameters to other parameters and reduces activation memory by scaled activation pruning.
G.3 Extra Information
The experiments are simulated via multi-process on an Nvidia RTX 5880 GPU, an Intel Core i9 CPU with 48 GB of memory.
G.4 Communication and Computational Cost